# GreenScan: Towards large-scale monitoring the health of urban trees using mobile sensing

### 22 Dec 2023 v1

@@ Abstract
Healthy urban greenery is a fundamental asset to mitigate climate change phenomenons such as extreme heat and air pollution. However, urban trees are often affected by abiotic and biotic stressors that hamper their functionality, and whenever not timely managed, even their survival. While current visual or instrumented inspection techniques can help take effective measures, they often require a high amount of human labor, making frequent assessments infeasible at a city-wide scale. In this paper, we present GreenScan, a ground-based sensing system designed to provide health assessment of urban trees at high spatio-temporal resolutions, with low costs. The system utilises thermal and multi-spectral imaging sensors fused using custom computer vision models in order to estimate two tree health indexes. The evaluation of the system was performed through data collection experiments in Cambridge, USA. Overall, this approach illustrates the potential of autonomous mobile ground-based tree health monitoring on city-wide scales at high temporal resolutions with low-costs.
I Introduction
@@
Urban greenery improves cities’ resilience to climate change. Nowadays, protecting, managing, and restoring ecosystems is fundamental for climate-resilient development, given the multiple risks posed to humanity and nature by global warming and climate change as per the latest UN-IPCC Report [1]. In urban cities, the tree canopies and vegetation provide a wide range of ecosystem services such as air filtering, carbon sequestration, reduced energy consumption, and decreased local temperatures [2, 3]. However, urban trees are experiencing an ample amount of abiotic and biotic stressors that are exacerbated due to climate change [4, 5, 6]. As a result, their functionality, productivity, and survival are of increasing concern [7]. Trees with poor health cannot provide most of their beneficial ecosystem services [8, 9]. For instance, trees with low transpiration rates do not cool the environment sufficiently and trees with low growth rates have a reduced shading effect. By 2050, it is expected that about two-thirds of urban tree species worldwide will fail to provide the desired climate-positive benefits [10].
Nowadays, the health of trees can be inspected by arborists with good quality results, but at high costs due to the amount of human labor involved [11]. Therefore, leading to an assessment that has a low spatial and temporal resolution. In recent years, a few technology-assisted methods have been developed to complement manual inspections. Satellite-based imaging can cover large areas although at a low spatial granularity [12], with data quality depending on the availability of clear skies [12] and low-temporal resolution because satellites revisit the same spot only every few days. Airborne sensing using Unmanned Aerial Vehicles (UAVs) or airplanes leads to an increased spatial granularity [12]; yet it involves high operational costs and may not be suitable in highly urbanized environments due to aviation authority regulations. Further, both airborne sensing and satellite imagery can only capture the overhead view of the urban canopy. As a result, vegetation elements such as green walls, short trees, or shrubs present under the tree canopy may be missed [13] or misinterpreted [14]. These challenges make it economically infeasible for governments and municipalities to perform frequent inspections of urban greenery at a large-scale. As a result, adverse health conditions are often discovered only after severe damage and, sometimes, beyond the point of no return. Further, given the lack of tree health data, intricate relationships of urban trees with other micro-scale ecosystem services is not well understood from an urban planning perspective. For instance, inappropriate placement of trees in outdoor environments can be detrimental as they can serve to reduce and trap air pollutants while also hindering anthropogenic heat [15].
@@
Recently, a number of projects have investigated the use of low-cost technological alternatives to monitor the environmental parameters in cities. For instance, applying Artificial Intelligence (AI) based methods on Google Street View (GSV) images to detect the presence of trees [16, 17], or using drive-by sensing strategies to measure air pollution and other urban morphology parameters [18]. In the case of drive-by sensing (retrofitting vehicles with sensors), it was demonstrated that just ten random taxis could capture data over one-third of streets in Manhattan (New York City) in a single day [19]. Drive-by sensing approaches enable collecting data at both higher temporal and spatial resolutions compared to manual inspection approaches, embedded sensors, or satellite-based sensing. Supplementary, citizen science-based approaches [20] have also been successful in measuring urban environmental parameters such as air pollution [21]. All these methods are set within the domain of opportunistic sensing and are aimed at developing platforms that can be deployed and operated without the need of an expensive or a dedicated infrastructure.
@@
Following on this trend and the critical need for protecting and managing urban ecosystems, in this work, we investigate how to measure the health of urban trees, terrestrially, at high spatial and temporal resolutions with low costs for wide-accessibility. The system described in this paper namely, GreenScan, measures the health of urban trees on a city-wide scale using complementary indexes that represents a tree’s photosynthetic capacity and a tree’s water stress levels. The system fuses data from low-cost thermal and multispectral imaging sensors using custom computer vision models to generate two popular health indexes namely NDVI (Normalised Difference Vegetation Index) and CTD (Canopy Temperature Depression). NDVI is based on the optical properties of chlorophyll and cell structure in the leaves [22], while CTD is physically related to the overall water stress of the trees [23]. A GreenScan system is suitable for deployment in a citizen-science paradigm by being carried by pedestrians and in a drive-by sensing paradigm by mounting on urban vehicles such as taxis and garbage trucks. Thus, allowing large-scale monitoring the health of trees at high spatial and temporal resolutions for cities and municipalities around the world.
@@
In this paper, we first give a brief overview of the existing state-of-the-art technology supported tools and methods to monitory health of urban greenery and identify the relevant research gaps. We use this to motivate the design of our system and define the architecture of GreenScan. A concrete implementation of both the hardware and software components is also described. We evaluate the performance of the system in measuring the health of trees while also discussing the performance of the efficiency optimised computer vision model. We conclude by identifying the immediate future work enabled through large scale deployment of GreenScan and illustrate the overall potential of autonomous ground-based tree health monitoring.
II Related Work
@@
A comparison of the surveyed methods currently used for tree health monitoring in terms of working mechanism, cost, and quality of assessment is shown in Table I.
Approach Working Mechanism Quality of Assessment Cost*
@@
Handheld imaging Hyper/multi-spectral imaging
Thermal imaging
LiDAR
@@
Remote sensing (hyper/multi-spectral, thermal imaging, LiDAR)
@@
TABLE I: A comparison of sensing approaches along with suitable deployment strategies to analyse tree health
(* refers to relative cost where \$ is the lowest cost and \$\$\$\$ is the highest cost for large-scale evaluation of multiple trees )
@@
Manual inspection involves the work of arborists (a tree doctor) inspecting each tree visually, often with the aid of tools like borers (to extract a wooden core sample from the tree for laboratory analysis) or resistographs (to measure the electrical resistance of the trunk). These methods usually provide high-quality assessment, but they are time-consuming due to the amount of human labor involved to perform a tree-by-tree assessment. Further, although effective, methods that require drilling and penetration in the living wood may create an entry path for pathogens or may alter the structural integrity of a tree. For a review on these methods, see [11].
@@
Embedded sensing involves the deployment of either physical or chemical sensors, which are attached on or near the tree of interest. The physical property under scrutiny can vary, from the detection of sudden vibrations to the measurement of water uptake and transpiration. Usually, this family of methods generates data at high temporal resolutions with little or no human supervision required; yet at the cost of installing and maintaining one or more sensors per tree. For a review on these methods, see [24].
@@
Imaging-based methods involve the use of optical sensors such as thermal imaging sensors, multispectral or hyperspectral imaging sensors or LiDAR (Light Detection and Ranging). Thermal imaging is based on IR (InfraRed) radiation emitted from biological materials and it is mainly used to (i) measure cavities in the living wood, (ii) detect infections caused by insects and bacteria, and (iii) calculate water stress levels by measuring the temperature of the leaves in the canopy. On the other hand, multi/hyper-spectral imaging captures various bands in the electromagnetic spectrum, usually near-infrared and parts of visible spectrum. This captured data is utilised to calculate various vegetation indexes, the most popular being Normalized Difference Vegetation Index (NDVI) which requires the visible red radiation and the Near InfraRed (NIR) radiation reflected by the vegetation. High NDVI values are a proxy for good health or healthy photosynthetic capacity. Low NDVI values are a proxy for overall poor health, the presence of stress or parasites, or the absence of vegetation elements [25]. Light Detection and Ranging (LiDAR) sensors, can be utilised to measure geometrical parameters such as the leaves surrounding a branch, trunk diameter, etc. and to estimate the Leaf Area Index (LAI) [26]. These imaging based approaches are non-invasive since they don’t require any physical contact with the tree, in comparison to embedded sensing and manual inspection approaches.
@@
Recently, street-view based methods based on visible RGB (Red, Green, Blue) images such as [27, 28, 29, 30], involving the use of google-street view images are emerging. While these approaches are cost-effective and scalable, they are only able to quantify the extent of urban greenery at a terrestrial level rather than its health.
@@
Finally, remote sensing utilising imaging-based sensors deployed on satellites, and Unmanned Aerial Vehicles (UAVs) or airplanes enable high spatial coverage. However, satellites have a low temporal resolution due to infrequent re-visit time over the same spot and the data quality is dependent on the availability of clear skies [12]. Data collection using UAVs and airplanes involves high operational costs and is unsuitable for highly urbanised environments due to aviation regulations. Most importantly, both airborne sensing and satellite imagery can only capture an overhead view of urban tree canopies. As a result, lower vegetation elements such as green walls, short trees, or shrubs are often missed or misinterpreted [13]. For a review on these methods, see [31].
II-A Research gaps and influence on design
Comparing different approaches, it emerges that manual inspection methods are infeasible to be deployed on city-side scales due to the high costs and the amount of human labour involved. Further, a subset of these methods involve the use of invasive tools. While deploying imaging-based sensors on satellites and drones can cover large areas; the resulting images have pixels with low spatial granularity, low temporal resolutions due to revisit time [12], and only provide an overhead view of the vegetation. Further, the presence of background materials such as grass or shrubs and clouds in the atmosphere affects the reliability of observations [32].
In contrast, ground-based sensing approaches combined with imaging-based sensors can look at vegetation elements in a holistic manner. For instance, utilising thermal cameras to analyze health attributes such as water stress index and cavities appears theoretically suitable and promising. Hyperpectral/multispectral imaging can also generate a number of vegetation health indexes, the most popular of which is NDVI. Contradictory studies were observed on the usage of LiDAR with some works such as [33] claiming no increase in health classification performance with its addition. Overall, except for a few works utilising imaging-based methods for terrestrial monitoring such as [34, 35], the remaining surveyed works required the use of manual handheld sensor data collection along with manual data processing. Thus, incorporating substantial human labor due to the absence of automated processes, thereby constraining the attainable spatial and temporal resolutions. With the advancements in computer vision and deep learning, in GreenScan, we utilise both multi-spectral and thermal imaging sensors to autonomously measure tree health indexes from ground-level.
II-B Comparison with comparable earlier works
Our work is unique in the sense that it is suitable to be deployed for large-scale monitoring of trees without any human intervention. Earlier works in the field which measured tree health terrestrially and used low cost sensors are shown concisely in Table II along with the research gaps in them and how our works fills this research gap. The two utmost comparable works; [35] requires manual human analysis of images and [12] outputs only raw values without ground truth validation combined with controlled upward facing deployment respectively. Our work fulfills this research gap by being completely autonomous and suitable to deployed for opportunistic sensing along with validated ground truth data provided by a municipality.
@@ II-B Comparison with comparable earlier works
172 trees (only system output as no comparison)
@@
2 trees in controlled lab environment
@@
Our work Yes Mobile (Cars) and Citizen Science Yes
@@
TABLE II: A concise comparison of our work with comparable ealier works in the field.
III Methodology
III-A System Architecture
@@
The block diagram of the entire GreenScan system architecture is shown in Figure 1. Herewith, we give a brief description of all the other major modules, as shown in the architecture diagram. The first five modules are related to hardware, while the remaining four modules are related to software.
All the hardware components were encased in a 3D printed case as shown in Figure 2. The case was designed such that it is suitable to be attachable to moving vehicles without any alterations using magnets, as shown in Figure 3.
Figure 2: All hardware components encased with the 3D printed case
Figure 3: Concept casing: (a) Concept casing for the system with magnets; (b) The system attached to the top of a car; (c) A closeup view of the system attached to the roof of a car
III-A1 Hardware Modules
Below, we first provide the generic description of each hardware module followed by the concrete implementation of the same in our prototype GreenScan system.
@@ III-A1 Hardware Modules
1. Thermal Imaging Sensor: A thermal imaging sensor with radiometric calibration (to measure true temperature) is attached to the central single board computer and captures long-wave infrared images normalised to a suitable temperature range with low pixel resolution. A narrow temperature range is preferred to decrease the effect of non-linear noise across the sensor and the low cost thermal imaging sensors are constrained in terms of resolution. This long-wave infrared imaging data is used for generation of CTD (Canopy Temperature Depression). In future, this thermal imaging data can also be used for detection of cavities in the trunk of the trees as proposed in [40]. However, this was skipped in our prototype mainly due to the low-resolution of the thermal imaging sensor.
@@
For concrete implementation, we used FLIR Lepton 3.5 (spectrum: longwave-infrared @8 µm - 14 µm) attached to an OpenMV cam H7 using a FLIR Lepton adapter module. This captured thermal images with a pixel resolution of $160\times 120$ which are normalised to a suitable temperature range ($-10^{\circ}$ to $40^{\circ}$ C). This temperature range was chosen based on the lowest and highest temperature (±$\pm 10^{\circ}$) of trees found during the data collection experiments (in Section IV). The OpenMV cam H7 communicates with Raspberry Pi via remote procedure call (RPC) over USB.
@@
2. Multispectral imaging sensor: A multispectral imaging sensor is attached to the central single board computer and captures Red, Green and Near-infrared (RGN) imaging data with high pixel resolution. The near-infrared and red imaging data is used for generation of the NDVI (Normalised Difference Vegetation Index). Further, this high resolution imaging data is used for segmentation of the tree canopy from the images using the custom deep learning model as described in the Image Segmentation Module.
@@
For implementation, MAPIR Survey 3W (spectrum: red@660nm, green@550nm, near-infrared@850nm) was attached to the Raspberry Pi over USB and captured RGN imaging data with a pixel resolution of $4000\times 3000$. To control the MAPIR Survey 3W for triggering the capture and transfer of images, pulse width modulation (PWM) signals over the micro-HDMI port of MAPIR Survey 3W are used.
@@
3. GNSS Receiver: A Global Navigation Satellite System (GNSS) receiver with support for GPS, GLONASS, and Galileo is used to find the current location of the system and geo-tag all the images of the trees captured.
@@
For our prototype, the RGN images captured were geo-tagged using the standard GPS adapter available for MAPIR Survey 3W.
@@
4. Single Board Computer with/without Edge TPU: A single board computer without/with onboard edge Tensor Processing Unit (TPU) co-processor/USB edge TPU accelerator (such as USB TPU accelerator from coral.ai) acts as the central brain of the system integrating all the hardware and software components. The edge TPU allows to speed up deep learning operations, improving the images processed per second without sending any data to the cloud.
@@
Raspberry Pi 3 was utilised as the single board computer in the GreenScan system running the software modules (see Section III-A2).

### 06 Apr 2024 v2

# GreenScan: Towards large-scale terrestrial monitoring the health of urban trees using mobile sensing

@@ Abstract
Healthy urban greenery is a fundamental asset to mitigate climate change phenomena such as extreme heat and air pollution. However, urban trees are often affected by abiotic and biotic stressors that hamper their functionality, and whenever not timely managed, even their survival. While the current greenery inspection techniques can help in taking effective measures, they often require a high amount of human labor, making frequent assessments infeasible at city-wide scales. In this paper, we present GreenScan, a ground-based sensing system designed to provide health assessments of urban trees at high spatio-temporal resolutions, with low costs. The system utilises thermal and multi-spectral imaging sensors fused using a custom computer vision model in order to estimate two tree health indexes. The evaluation of the system was performed through data collection experiments in Cambridge, USA. Overall, this work illustrates a novel approach for autonomous mobile ground-based tree health monitoring on city-wide scales at high temporal resolutions with low-costs.
1 Introduction
[…]
@@ 1 Introduction
Urban greenery improves the resilience of cities to climate change. Nowadays, protecting, managing, and restoring greenery ecosystems is fundamental for climate-resilient development, given the multiple risks posed to humanity and nature by global warming and climate change as per the latest UN-IPCC Report [1]. In cities, tree canopies and vegetation provide a wide range of ecosystem services such as air filtering, carbon sequestration, reduced energy consumption, increased biodiversity and decreased local temperatures [2, 3]. However, urban trees are experiencing an ample amount of abiotic stressors (e.g. soil salinity, heat waves) and biotic stressors (caused by living agents such as insects and bacteria) that are exacerbated due to climate change [4, 5, 6]. As a result, their functionality, productivity, and survival are of increasing concern [7]. Trees with poor health cannot provide most of their beneficial ecosystem services [8, 9]. For instance, trees with low transpiration rates do not cool the environment sufficiently and trees with low growth rates have a reduced shading effect. By 2050, it is expected that about two-thirds of urban tree species worldwide will fail to provide the desired climate-positive benefits [10].
The practice of measuring and monitoring urban trees began over a century ago [11]. Today, the health of trees can be monitored using manual inspection by arborists with good quality results [12]; yet the high labor cost lead leads to assessments performed infrequently at very low temporal resolutions such as once every 3-5 years. Technology-assisted monitoring methods can complement manual inspections [13]. However, these methods are impeded by variable data quality, low spatial granularity (remote sensing), or high operational costs (airborne sensing) [14]. Further, most of these methods are unable to quantify the vegetation elements below the tree canopy such as green walls, short trees, or shrubs [15, 16]. All these challenges lead to the lack of urban tree health data in cities and appropriate urban forest management. For instance, adverse health conditions in trees being discovered only after severe damage is already inflicted. Further, from an urban planning perspective, intricate relationships of urban trees with other micro-scale ecosystem services such as air quality improvements and benefits to public health are difficult to quantify. For instance, inappropriate placement of trees in outdoor environments can be detrimental as they can serve to trap air pollutants [17].
@@
Recently, several projects have investigated developing novel alternatives for environmental sensing. For instance, applying AI (Artificial Intelligence) based methods on GSV (Google Street View) images to detect the presence of trees [18, 19], or using drive-by strategies to measure air pollution [20, 21] in a cost efficient way. For instance, it was demonstrated that just ten random taxis could capture data over one-third of streets in Manhattan (New York City) in a single day using drive-by sensing [22]. Additionally, citizen science-based approaches [23] have also been successful in measuring urban environmental parameters [24]. All these methods are set within the domain of opportunistic sensing and are aimed at developing platforms that can be deployed and operated without the need of an expensive or a dedicated infrastructure. Thus, allowing democratic access to even under-resourced cities that are affected by climate change in a disproportionate manner [25].
@@
Following on this trend and the critical need for protecting and managing urban forestry, in this work, we develop a novel system, named GreenScan, which measures the health of urban trees on city-wide scales from ground level (terrestrially). The system fuses high-quality data from low-cost thermal and multispectral imaging sensors using custom computer vision models to generate two complementary tree health indexes namely NDVI (Normalised Difference Vegetation Index) and CTD (Canopy Temperature Depression) which respectively indicate the photosynthetic capacity and water stress levels of a tree. GreenScan was designed both for deployment in citizen-science paradigms by being carried by pedestrians, or in drive-by sensing approaches by being mounted on urban vehicles such as taxis and garbage trucks. Thus, enabling terrestrial urban tree health measurements with high spatial and temporal resolutions at low costs for cities and municipalities around the world.
@@
In this paper, we first give a brief overview of the state-of-the-art technology tools and methods to monitor the health of urban greenery with focus on low-cost solutions. We present the design of GreenScan and describe the implementation of the hardware and software components. We evaluate the system with forty urban trees in uncontrolled outdoor environments and analyse the performance of the system. Finally, we conclude by identifying the immediate future research that can be enabled through large scale deployment of GreenScan while discerning the limitations of this work.
2 Related Work
Currently, the health of trees is monitored through manual inspection by human experts, remote and airborne sensing through satellites or UAVs, direct installation of embedded sensors on/near the tree, handheld imaging based sensing or opportunistic sensing using street view imaging [13]. A comparison of these methods in terms of working mechanism, cost, and quality of assessment is shown concisely in Table 1.
@@ 2 Related Work
Table 1: A comparison of sensing approaches along with the working mechanism, cost, and quality of assessment to analyse tree health
(* refers to relative cost where \$ is the lowest cost and \$\$\$\$ is the highest cost for large-scale evaluation of multiple trees based on the scale in [12])
Approach Principle Quality of Assessment Cost*
@@
Handheld imaging Multi/Hyper-spectral imaging Thermal LiDAR
@@
Remote sensing (multi/hyper-spectral, thermal imaging, LiDAR)
@@
Manual inspection involves the work of arborists (human-experts) inspecting trees visually, often with the aid of tools such as borers (to extract a wooden core sample from the tree for laboratory analysis) or resistographs (to measure the electrical resistance of the trunk). These methods usually provide a high-quality assessment, but they are time-consuming due to the amount of human labor involved to perform a tree-by-tree assessment. Further, although effective, methods that require drilling and penetration in the living wood may create an entry path for pathogens or may alter the structural integrity of a tree. For a review on these methods, see [12].
@@
Embedded sensing involves the deployment of sensors in the bark of a tree (the outer wooden part of a tree) or in the soil. These sensors can rely on physical, chemical or electrical phenomenon to detect the presence of parasites, e.g. detecting sudden minimal bark vibrations produced by parasites locomotion and feeding [26] as well as water uptake and transpiration, e.g. measuring electrical impedance using a pair of electrodes placed in the trunk at opposite positions [27]. These methods generate data at high temporal resolutions with little or no human supervision required; yet at the cost of installing and maintaining one or more sensors per tree. For a review on these methods, see [28].
@@
Imaging-based methods involve the use of optical sensors such as thermal imaging sensors, HMI (hyperspectral or multispectral imaging) sensors or LiDAR (Light Detection and Ranging). Thermal imaging is based on IR (InfraRed) radiation emitted from materials and it is mainly used to (i) measure cavities and physical damages in the living wood [29, 30], (ii) detect infections caused by insects and bacteria [31, 32], and (iii) calculate water stress levels by measuring the temperature of the leaves in the canopy [14]. On the other hand, HMI sensors capture various bands in the electromagnetic spectrum, usually near-infrared and parts of the visible spectrum. This captured data is utilised to calculate various vegetation indexes, the most popular being NDVI (Normalized Difference Vegetation Index). HMI sensors are often used for remote sensing applications [33], although static sensors also exist [34]. Calibration methods are critical to achieve quality results [35]. LiDAR (Light Detection and Ranging) sensors, can be utilised to measure geometrical parameters such as the leaves surrounding a branch, trunk diameter, etc. and to estimate the LAI (Leaf Area Index) [36]. However, contradictory studies have been observed on the usage of LiDAR with some works such as [37] claiming no increase in health classification performance with its addition. Usually, LiDAR and HMI sensor approaches are often deployed in tandem in both airborne [36] and ground-based [38] approaches.
@@
Recently, street-view based methods based on RGB (Red, Green, Blue) images usually involving the use of google-street view images have became popular. These methods are utilised either to quantify the presence of urban greenery [18, 39], catalog species [40], and shading effects [41]. While these approaches are cost-effective and scalable, they are only able to quantify the extent of urban greenery at a terrestrial level rather than its health.
@@
When imaging-based sensors are deployed on satellites, airplanes or UAVs (Unmanned Aerial Vehicles), high spatial coverage can be achieved. However, satellites have a low temporal resolution due to infrequent re-visit time and data quality being dependent on the availability of clear skies [14]. Data collection using UAVs and airplanes involves high operational costs and is unsuitable for highly urbanised environments due to aviation regulations. Most importantly, both airborne sensing and satellite imagery can only capture an overhead view of urban tree canopies. As a result, lower vegetation elements such as green walls, short trees, or shrubs are often missed or misinterpreted [15].
For a systematic review of the technological methods and tools for greenery health monitoring, see [13].
2.1 Research gaps and influence on design
We seek to provide a scalable system that provides high quality data with low-costs. Comparing the different approaches in Table 1, it emerges that ground-based (terrestrial) sensing approaches combined with imaging-based methods can look at vegetation elements in a holistic manner with high quality data gathered either through drive-by sensing or citizen science paradigms. Additionally, the advances in deep learning based computer vision models for imaging data in the past decade enables the development of a system that is broad in scope. Narrowing this down, the past studies measuring tree health from ground level and utilising low-cost imaging sensors are also shown concisely in Table 2. These studies are limited by requiring manual analysis of images by humans [42], outputting only raw data without ground truth validation [14, 43] or requiring controlled system deployment and operation [14, 44, 45].
Our work builds upon all these insights. Hence, in GreenScan, we utilise HMI imaging and thermal imaging sensors to autonomously measure two health indexes namely NDVI and CTD from ground level. GreenScan is designed to be completely autonomous (by utilising computer vision model based on deep learning), is suitable to be deployed in non-controlled environments and the early results are compared with a ground truth dataset provided by a municipality. While data from HMI and thermal imaging can be used to generate a number of health indices, we carefully chose NDVI and CTD indices after explicit considerations. Our choices are driven by 1) NDVI remains one of the most important and popular indices used in the domain [35] and the ground truth dataset provided by the municipality contains remote NDVI for a fair scientific comparison 2) CTD is one of the relatively simpler metrics for assessing properties of a tree such as water consumption and its resilience to drought and heat stress events [46], and it uses different wavelengths than NDVI (thermal imaging sensors instead of HMI sensors). In turn, generating two complementary health parameters for urban trees.
Table 2: A concise comparison of our work with earlier works in the field measuring tree health terrestrially
@@ 2.1 Research gaps and influence on design
172 trees (only raw system output with no comparison)
@@
2 trees (in controlled lab environment)
@@
This work Yes Mobile (Cars) and Citizen Science Yes
@@
3 Methodology
The GreenScan system integrates low-cost thermal and multispectral imaging sensors which are attached to a single board computer. The system processes the imaging data generated by these sensors using a custom computer vision model to generate the two tree health indexes namely CTD and NDVI. All these components were encased in a 3D printed case as shown in Figure 1(a). The case was designed such that it is suitable to be attachable to moving vehicles without any alterations using magnets, as shown in Figures 1(b), 1(c) and 1(d). In this section, we aim to explain all the major modules of GreenScan.
3.1 System Architecture
@@
The block diagram of the entire GreenScan system architecture is shown in Figure 1. The first five modules are related to hardware, while the remaining four modules are related to software.
(a) All hardware components encased with the 3D printed case. Dimensions in inches: 7.32” X 2.50” X 1.74”
(b) Concept casing with magnets
(c) The system attached on the top of a car
(d) Closeup view of the system on the roof of a car
Figure 2: A visualization of the current GreenScan system and the concept casing
3.1.1 Hardware Modules
In the following, we first provide the generic description of each hardware module followed by the concrete implementation of the same in the GreenScan system.
@@ 3.1.1 Hardware Modules
1. Thermal Imaging Sensor: A thermal imaging sensor with radiometric calibration (to measure true temperature) is attached to the central single board computer and captures long-wave infrared images normalised to a suitable temperature range with low pixel resolution. A narrow temperature range is preferred to decrease the effect of non-linear noise across the sensor as the low cost thermal imaging sensors are constrained in terms of resolution. This long-wave infrared imaging data is used for generation of CTD (Canopy Temperature Depression) which indicates the water stress levels of a tree.
@@
For concrete implementation, we used FLIR Lepton 3.5 (spectrum: longwave-infrared @8 µm - 14 µm) attached to an OpenMV cam H7 using a FLIR Lepton adapter module. This captured thermal images with a pixel resolution of $160\times 120$ which are normalised to a suitable temperature range ($-10^{\circ}$ to $40^{\circ}$ C). This temperature range was chosen based on the lowest and highest temperature (±$\pm 10^{\circ}$) of trees found during the data collection experiments (in Section 4). The OpenMV cam H7 communicates with Raspberry Pi via RPC (remote procedure call) over USB.
@@
2. Multispectral imaging sensor: A multispectral imaging sensor is attached to the central single board computer and captures RGN (Red, Green and Near-infrared) imaging data with high pixel resolution. The near-infrared and red imaging data is used for generation of the NDVI (Normalised Difference Vegetation Index). Further, this high resolution imaging data is used for segmentation of the tree canopy from the images using the custom deep learning model as described in the Image Segmentation Module.
@@
For implementation, MAPIR Survey 3W (spectrum: red@660nm, green@550nm, near-infrared@850nm) was attached to the Raspberry Pi over USB and captured RGN imaging data with a pixel resolution of $4000\times 3000$. To control the MAPIR Survey 3W for triggering the capture and transfer of images, PWM (pulse width modulation) signals over the micro-HDMI port of MAPIR Survey 3W are utilised.
@@
3. GNSS Receiver: A GNSS (Global Navigation Satellite System) receiver with support for GPS, GLONASS, and Galileo is used to find the current location of the system and geo-tag all the images of the trees captured.
@@
For this prototype, the RGN images captured were geo-tagged using the standard GPS adapter available for MAPIR Survey 3W.
@@
4. Single Board Computer with/without Edge TPU: A single board computer without/with onboard edge TPU (Tensor Processing Unit) or USB edge TPU accelerator (such as USB TPU accelerator from coral.ai) acts as the central brain of the system integrating all the hardware and software components. The edge TPU allows to speed up deep learning operations, improving the images processed per second without sending any data to the cloud.
@@
Raspberry Pi 3 was utilised as the single board computer in the GreenScan system running the software modules (see Section 3.1.2).